# Sensitivity and topology of exceptional rings in nonlinear non-Hermitian planar optical microcavities

### v1

@@ V Topology of Exceptional Rings
Figure 8: Topological structure of nonlinear exceptional rings in 4D parameter space. (a) With Zeeman splitting, a single linear exceptional ring at $\Delta_{\mathrm{B}}=0$, in the nonlinear regime expands into a quasi-torus of exceptional nexuses (colored) connected by surfaces of exceptional arcs (black). (b) Rashba-Dresselhaus splitting in $k_{\parallel}$-direction leads to a tilt of the quasi-torus along $k_{\parallel}$ (shown for $\alpha_{\mathrm{RD}}=0.1~{}\mathrm{meV}$).
@@
We use the solutions of the nonlinear eigenvalue problem (3) with finite detuning to investigate topological features by applying Rashba-Dresselhaus and Zeeman splitting to the system. In contrast to the linear ER, the elliptic umbilic allows the formation of new ER constellations for finite detuning $\Delta$. In Fig. 7 the complex-valued energy spectrum and central cuts are displayed for a finite Zeeman splitting of $\Delta_{\mathrm{B}}=12.5~{}\mathrm{\upmu eV}$. The inset in Fig. 7(b) shows the umbilic cut at $\alpha=0.1~{}\mathrm{meV\upmu m}^{2}$ in turquoise together with the chosen Zeeman splitting in orange. The intersections mark the visible ERs for the given setup. Here, two rings are formed by points on the exceptional arcs. Notably, reversing the direction of the Zeeman-induced energy shift inverts the polarization in the complex-valued energy spectrum. We note that similar energy spectra can be achieved for stronger Zeeman splitting if the nonlinearity is increased simultanously (to increase the size of the umbilic in the regime where its structure is valid). A relatively weak nonlinearity was selected to limit the blueshift, ensuring the resulting structures remain within an easily visualizable energy range. These findings highlight, that while they both feature identical $1^{\mathrm{st}}$ Chern numbers, the neighborhood topology of linear and nonlinear ERs is drastically different, as detuning is introduced.
@@ VI Conclusion
We have demonstrated that exceptional rings (ERs) naturally emerge inside the in-plane dispersion of planar optical resonators with TE-TM splitting and circular dichroism (leading to different losses in the oppositely circularly polarized resonator modes). These ERs can even be recovered in $k$-space-anisotropic systems, where adjusting mode detuning can be used to relocalize them in higher-dimensional parameter space. Introducing a Kerr-type medium (or saturable gain) splits the linear ER into mutliple ERs forming an elliptic umbilic in nonlinear parameter space. Here, $3^{\mathrm{rd}}$-order exceptional nexus rings and $2^{\mathrm{nd}}$-order exceptional arc rings form concentrically in $k$-space. We performed a rigorous analysis on the perturbation response of the nonlinear ERs, showing enhanced signal enhancement and diverse perturbation-response functions for ERs at different positions on the umbilic singularity structure. We show that a linear ER in higher-dimensional parameter space splits into a complex quasi-toroidal structure of ERs, making nonlinear ERs a promissing candidate for mode-switching applications. Since a wide range of nonlinear non-Hermitian Hamiltonians feature exceptional points or contours that are organization point of an elliptic umbilic [52], our results on topology and perturbation response can be applied to a large group of physical systems hosting exceptional points that are subject to real (or imaginary) valued smooth nonlinearities. Examples are Kerr-type nonlinearity or saturable gain, but also higher-order nonlinearities and nonlinearities of different functional form.
@@ Acknowledgements.
The authors gratefully acknowledge financial support for the Paderborn groups, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Grant No. 467358803 and the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078), and for the Arizona group from the US National Science Foundation (NSF) under Grant No. DMR-1839570. The authors gratefully acknowledge the computing time provided to them on the high-performance computer Noctua 2 at the NHR Center PC2 under project hpc-prf-hermep. LA and JS acknowledge support through the QuantERA project QuCABOoSE. We thank Jan Wiersig and Julius Kullig for fruitful discussions, specifically on the aspects of nonlinearity and signal enhancement.
@@ Appendix A Exceptional ring generation and control by off-resonant pumping
Here, the field is coupled to an incoherent excitation reservoir $n_{\pm}$ driven by continous-wave pump $P_{\pm}$. The excitation reservoir exhibits a loss rate $\gamma_{\mathrm{r}}$ and its scattering into the field is determined by the stimulated scattering rate $r$. The loss term now holds $\Gamma_{\pm}=\frac{\hbar}{2}[rn_{\pm}-\gamma_{\pm}]$. The pump is below threshold to avoid condensation, however, in prinicple the eigenvalue problem (1) shows an exceptional ring in the gain regime as well. For the calculations with the extended model in equations (8,9) we find the results (not shown) to be in perfect agreement with the results where the difference in loss rate was introduced by hand in Fig. 2(a) above. Note that to realize the exceptional ring in this case the reservoir-induced blueshift $g_{r}n_{\pm}$ resulting from polariton-exciton interaction has to be compensated (for example by Zeeman splitting) to preserve mode degeneracy. The results displayed in Figure 2 can be reproduced with the polarized reservoir as source for circular dichroism.
@@ Appendix B Analytical solution of the nonlinear eigenvalue problem at zero detuning
In this section we elaborate on the method used to calculate the analytical solutions of the nonlinear eigenvalue problem (3) without detuning, inspired by the separability eigenvalue problem [81]. This is due to the fact that nonlinearities of vector-valued quantities can be expressed in terms of tensor products, such as $|a\rangle^{\otimes 2}=|a\rangle\otimes|a\rangle$ representing a square of the vector $|a\rangle$ in the tensor algebra [98]. First we make the valid assumption $\Gamma_{1},\Gamma_{2},g_{x},g_{x},k_{x},k_{y},\Delta_{\mathrm{LT}}\in\mathbb{R}$. To this end we rewrite the eigenvalue problem:
$\begin{bmatrix}H&-\Delta_{\mathrm{LT}}|\textbf{k}|^{2}\\ -\Delta_{\mathrm{LT}}|\textbf{k}|^{2}&-H\end{bmatrix}\begin{bmatrix}\psi_{+}\\ \psi_{-}^{\prime}\end{bmatrix}=\mu^{\prime}\begin{bmatrix}\psi_{+}\\ \psi_{-}^{\prime}\end{bmatrix}$ (10)
with $H=i\frac{\Gamma_{1}-\Gamma_{2}}{2}+\frac{g_{c}-g_{x}}{2}(|\psi_{+}|^{2}-|\psi_{-}^{\prime}|^{2})$, $\mu^{\prime}=\mu-\mu_{0}$, $\mu_{0}=\frac{\hbar^{2}\textbf{k}^{2}}{2m}+\frac{i(\Gamma_{1}+\Gamma_{2})+g_{c}+g_{x}}{2}$, $\psi_{-}^{\prime}=\psi_{-}e^{-2\mathrm{i}\varphi}$ and $\varphi=\mathrm{arg}(k_{x}+\mathrm{i}k_{y})$. Then, we start deriving the operator for the separability eigenvalue equation [99, 81]
$\hat{L}_{a}|a\rangle=\mu^{\prime}|a\rangle.$ (11)
with partially reduced operator
$\hat{L}_{a}=(\langle a|\otimes\hat{1})\hat{L}(|a\rangle\otimes\hat{1})$ (12)
@@
where $\hat{1}$ denotes the identity and $\otimes$ the respective tensor product. Note that $\hat{L}_{a}$ can be also understood in terms of the computational basis $\{|i\rangle:i=0,1\}$. That is, the matrix elements of the partially reduced operator are $\langle i|\hat{L}_{a}|j\rangle=\langle a,i|\hat{L}|a,j\rangle$, with the common tensor-product notation $|a\rangle\otimes|b\rangle=|a,b\rangle$.
The general form of exchange-symmetric $\hat{L}$ expanded into Pauli matrices reads
@@
with $|a\rangle=[a_{0}~{}a_{1}]^{\mathrm{T}}$. This leads to the following nonlinear eigenvalue equation:
$=\mu^{\prime}\begin{bmatrix}a_{0}\\ a_{1}\end{bmatrix}$
which shows the same form as the rewritten eigenvalue problem in Eq. (10) and refers to the separability eigenvalue equation and orginates from entanglement theory [99, 81]. Using the solution ansatz
for the separability eigenvalue equation we can find the new equation
$0=\left(\sinh(\eta)\cos(\vartheta)-\frac{\alpha}{\beta}\frac{\sinh(\eta)}{\cosh(\eta)}\right)+i\left(\cosh(\eta)\sin(\vartheta)+\frac{\gamma}{\beta}\right),$ (19)
solving this equation yields the solutions of the rewritten separability equation:
In each of the two regions indicated in this equation we have two solutions, as can be seen from the $\pm$ in front of the square roots. However, there is a region where the two regions overlap, and in this overlap region we therefore have a total of four solutions. Here, the resulting eigenstates feature arbitrary phases. The calculated eigenvectors can then be inserted into the eigenvalue problem to determine the potentially complex eigenvalues of the nonlinear non-Hermitian system. The parameters $\alpha,\beta,\gamma$ can then be identified with the prefactors in equation (10). In the nonlinear regime, $\alpha\neq 0$, all four solutions of $w$ apply, leading to multistability. Note that the solutions are normalized to $|\psi_{+}|^{2}+|\psi_{-}|^{2}=N^{2}$, where $N>0$. For $N\neq 1$ the rewritten eigenvalue equation (10) can be rescaled by substituting $\tilde{a}_{0}=a_{0}/N$ and $\tilde{a}_{1}=a_{1}/N$, now obeying $|\tilde{a}_{0}|^{2}+|\tilde{a}_{1}|^{2}=1$. The left eigenvectors and eigenvalues are optained for $\hat{L}\mapsto\hat{L}^{\dagger}$, i.e., substitude $\gamma\mapsto-\gamma$ in the solutions.
Appendix C Generaliaztion of sensitivity and topological features to a large group of non-Hermitian Hamiltonians

### v2

@@ V Topology of Exceptional Rings
Figure 8: Topological structure of nonlinear exceptional rings in 4D parameter space. (a) With Zeeman splitting, a single linear exceptional ring at $\Delta_{\mathrm{B}}=0$, in the nonlinear regime expands into a quasi-torus of exceptional nexuses (colored) connected by surfaces of exceptional arcs (black). (b) Rashba-Dresselhaus splitting in $k_{\parallel}$-direction leads to a tilt of the quasi-torus along $k_{\parallel}$ (shown for $\alpha_{\mathrm{RD}}=0.1~\mathrm{meV}$). Simulation parameters: $\{\alpha,g_{\mathrm{c}},g_{\mathrm{x}}\}=\{0.1,0.2,0\}~\mathrm{meV\upmu m}^{2}$
@@
We use the solutions of the nonlinear eigenvalue problem (3) with finite detuning to investigate topological features by applying Rashba-Dresselhaus and Zeeman splitting to the system. In contrast to the linear ER, the elliptic umbilic allows the formation of new ER constellations for finite detuning $\Delta$. In Fig. 7 the complex-valued energy spectrum and central cuts are displayed for a finite Zeeman splitting of $\Delta_{\mathrm{B}}=12.5~\mathrm{\upmu eV}$. The inset in Fig. 7(b) shows the umbilic cut at $\alpha=0.1~\mathrm{meV\upmu m}^{2}$ in turquoise together with the chosen Zeeman splitting in orange. The intersections mark the visible ERs for the given setup. Here, two rings are formed by points on the exceptional arcs. Notably, reversing the direction of the Zeeman-induced energy shift inverts the polarization in the complex-valued energy spectrum. We note that similar energy spectra can be achieved for stronger Zeeman splitting if the nonlinearity is increased simultanously (to increase the size of the umbilic in the regime where its structure is valid). A relatively weak nonlinearity was selected to limit the blueshift, ensuring the resulting structures remain within an easily visualizable energy range. These findings highlight, that while they both feature identical quasi-Chern integrals, the neighborhood topology of linear and nonlinear ERs is drastically different, as detuning is introduced.
@@ VI Conclusion
We have demonstrated that exceptional rings (ERs) naturally emerge inside the in-plane dispersion of planar optical resonators with TE-TM splitting and circular dichroism (leading to different losses in the oppositely circularly polarized resonator modes). These ERs can even be recovered in $k$-space-anisotropic systems, where adjusting mode detuning can be used to relocalize them in higher-dimensional parameter space. Introducing a Kerr-type medium (or saturable gain) splits the linear ER into mutliple ERs forming an elliptic umbilic in nonlinear parameter space. Here, $3^{\mathrm{rd}}$-order exceptional nexus rings and $2^{\mathrm{nd}}$-order exceptional arc rings form concentrically in $k$-space. We performed a rigorous analysis on the perturbation response of the nonlinear ERs, showing enhanced signal enhancement and diverse perturbation-response functions for ERs at different positions on the umbilic singularity structure. We show that a linear ER in higher-dimensional parameter space splits into a complex quasi-toroidal structure of ERs, making nonlinear ERs a promising candidate for mode-switching applications. Since a wide range of nonlinear non-Hermitian Hamiltonians feature exceptional points or contours that are organization point of an elliptic umbilic [56], our results on topology and perturbation response can be applied to a large group of physical systems hosting exceptional points that are subject to real (or imaginary) valued smooth nonlinearities. Examples are Kerr-type nonlinearity or saturable gain, but also higher-order nonlinearities and nonlinearities of different functional form.
@@ Acknowledgements.
The authors gratefully acknowledge financial support for the Paderborn groups, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through Grant No. 467358803 and the transregional collaborative research center TRR 142 (Projects A04 and C10, Grant No. 231447078), and for the Arizona group from the US National Science Foundation (NSF) under Grant No. DMR-1839570. The authors gratefully acknowledge the computing time provided to them on the high-performance computer Noctua 2 at the NHR Center PC2 under project hpc-prf-hermep. LA and JS acknowledge support through the QuantERA project QuCABOoSE. We thank Jan Wiersig and Julius Kullig for fruitful discussions, specifically on the aspects of nonlinearity and signal enhancement. We further thank Anton Montag and Flore Kunst for a fruitful exchange on the role that symmetries of our system play in the linear regime.
@@ Appendix A Exceptional ring generation and control by off-resonant pumping
Here, the field is coupled to an incoherent excitation reservoir $n_{\pm}$ driven by continous-wave pump $P_{\pm}$. The excitation reservoir exhibits a loss rate $\gamma_{\mathrm{r}}$ and its scattering into the field is determined by the stimulated scattering rate $r$. The loss term now holds $\Gamma_{\pm}=\frac{\hbar}{2}[rn_{\pm}-\gamma_{\pm}]$. The pump is below threshold to avoid condensation, however, in prinicple the eigenvalue problem (1) shows an exceptional ring in the gain regime as well. For the calculations with the extended model in equations (10,11) we find the results (not shown) to be in good agreement with the results where the difference in loss rate was introduced by hand in Fig. 2(a) above, with minor differences in line profiles. Note that to realize the exceptional ring in this case the reservoir-induced blueshift $g_{r}n_{\pm}$ resulting from polariton-exciton interaction has to be compensated (for example by Zeeman splitting) to preserve mode degeneracy. The results displayed in Figure 2 can be reproduced with the polarized reservoir as source for circular dichroism.
@@ Appendix B Analytical solution of the nonlinear eigenvalue problem at zero detuning
In this section we solve the nonlinear eigenvalue problem (3) without detuning. First we make the valid assumption $\Gamma_{1},\Gamma_{2},g_{\mathrm{c}},g_{\mathrm{x}},k_{x},k_{y},\Delta_{\mathrm{LT}}\in\mathbb{R}$. For a better visualization of the eigenvalues, we substract the parabolic term, $\mu^{\prime}=\mu-\frac{\hbar^{2}\bm{k}^{2}}{2m}$, from the dispersion in the following. We write the eigenvalue problem as:
$\begin{bmatrix}H&\beta\\ \beta&-H\end{bmatrix}\begin{bmatrix}\psi_{+}\\ \psi_{-}^{\prime}\end{bmatrix}=(\mu^{\prime}-\mu_{0})\begin{bmatrix}\psi_{+}\\ \psi_{-}^{\prime}\end{bmatrix}$ (12)
with $H=i\frac{\Gamma_{1}-\Gamma_{2}}{2}+\frac{g_{c}-g_{x}}{2}(|\psi_{+}|^{2}-|\psi_{-}^{\prime}|^{2})$, $\mu_{0}=\frac{i(\Gamma_{1}+\Gamma_{2})+g_{c}+g_{x}}{2}$, $\psi_{-}^{\prime}=\psi_{-}e^{-2\mathrm{i}\varphi}$, $\varphi=\mathrm{arg}(k_{x}+\mathrm{i}k_{y})$, and $\beta=-\Delta_{\mathrm{LT}}|\bm{k}|^{2}$. After some rewriting this leads to the following nonlinear eigenvalue equation:
$=(\mu^{\prime}-\mu_{0})\begin{bmatrix}a_{0}\\ a_{1}\end{bmatrix}$
with $\gamma=\nicefrac{{|\Gamma_{1}-\Gamma_{2}|}}{{2}}$ and $\alpha=\nicefrac{{(g_{\mathrm{c}}-g_{\mathrm{x}})}}{{2}}$ as also introduced in the main part of the text and $a_{0}=\psi_{+}$ and $a_{1}=\psi_{-}^{\prime}$. We note that Eq (13) has the same form as the separability eigenvalue equation known in a more general context in entanglement theory [105, 85] as elaborated in Appendix C. Using the solution ansatz
we find the equation
$0$ $=\left(\sinh(\eta)\cos(\vartheta)-\frac{\alpha}{\beta}\frac{\sinh(\eta)}{\cosh(\eta)}\right)$ (16)
$+i\left(\cosh(\eta)\sin(\vartheta)+\frac{\gamma}{\beta}\right)\,.$
Solving this equation yields the solutions, determining the eigenvectors $|a\rangle$ with
In each of the two regions indicated in this equation we have two solutions, as can be seen from the $\pm$ in front of the square roots. However, there is a region where the two regions overlap, and in this overlap region we therefore have a total of four solutions. In his case, the eigenvalues can also be given in closed form. For non-zero $\beta$, we have
$\mu^{\prime}=\mu_{0}\pm\sqrt{\beta^{2}-\gamma^{2}}$ (17)
for $|\beta|\geq|\gamma|$ (region B and C in Fig. 4(a)), and
$\mu^{\prime}=\mu_{0}+\alpha\pm i\gamma\sqrt{1-\frac{\beta^{2}}{\alpha^{2}+\gamma^{2}}}$ (18)
for $|\beta|\leq\sqrt{\alpha^{2}+\gamma^{2}}$ (region A and B in Fig. 4(a)). This again shows that we have two solutions in region A and C, and four solutions in region B. Note, that the solution is not well-defined for $\beta=0$, i.e. at $k=0$ and $\gamma\neq 0$. The resulting eigenstates have arbitrary phases. In the nonlinear regime, $\alpha\neq 0$, all four solutions of $w$ apply, leading to multistability. Note that the solutions are normalized to $|\psi_{+}|^{2}+|\psi_{-}|^{2}=N^{2}$, where $N>0$. For $N\neq 1$ the rewritten eigenvalue equation (12) can be rescaled by substituting $\tilde{a}_{0}=a_{0}/N$ and $\tilde{a}_{1}=a_{1}/N$, now obeying $|\tilde{a}_{0}|^{2}+|\tilde{a}_{1}|^{2}=1$.
Appendix C Separability eigenvalue equations
Taking a more general view, the eigenvalue problem (12) can be cast into a form known from quantum information theory in the context of separability eigenvalue equations [85]. This is due to the fact that nonlinearities of vector-valued quantities can be expressed in terms of tensor products, such as $|a\rangle^{\otimes 2}=|a\rangle\otimes|a\rangle$ representing a square of the vector $|a\rangle$ in the tensor algebra [106].
In this section we derive the separability eigenvalue equation (13). We start deriving the operator for the separability eigenvalue equation [105, 85]
$\hat{L}_{a}|a\rangle=\lambda|a\rangle.$ (19)
with the partially reduced operator
[…]
@@ Appendix C Separability eigenvalue equations
where $\hat{1}$ denotes the identity and $\otimes$ the respective tensor product. Note that $\hat{L}_{a}$ can be also understood in terms of the computational basis $\{|i\rangle:i=0,1\}$. That is, the matrix elements of the partially reduced operator are $\langle i|\hat{L}_{a}|j\rangle=\langle a,i|\hat{L}|a,j\rangle$, with the common tensor-product notation $|a\rangle\otimes|b\rangle=|a,b\rangle$. The general form of exchange-symmetric $\hat{L}$ expanded into Pauli matrices reads
@@
with $|a\rangle=[a_{0}~a_{1}]^{\mathrm{T}}$. When substituted into Eq. (19), this results in an eigenvalue problem of the form of the rewritten eigenvalue problem (13).
Appendix D Generalization of sensitivity and topological features to a large group of non-Hermitian Hamiltonians
@@
Appendix E Topological properties for zero detuning
In this appendix, we evaluate topological properties of the ER in the $k_{x}-k_{y}$ plane, i.e. for the case of zero detuning. We do this first for the case without nonlinearity, and then extend the discussion to non-zero nonlinearity.
We define the closed-loop (sometimes called quantized) Berry phase of band $n$ calculated with left ($L$) or right ($R$) eigenvectors as
$\gamma^{ab}_{\mathrm{Berry,n}}=\oint_{\cal{L}}\bm{A}^{ab}_{n}(\bm{k})\cdot d\bm{k}$ (25)
[…]
Taking, for example, the upper band solution ($n=u$), we find the Berry phase for a closed loop in the $k_{x}-k_{y}$ plane with $\Delta=0$, after analytically diagonalizing Eq. (12),
$\gamma^{RR}_{\mathrm{Berry,u}}=2\pi\,{\rm mod}\,2\pi$ (26)
for $k_{\cal{L}}>k_{\mathrm{ER}}$, and
$\gamma^{RR}_{\mathrm{Berry,u}}=-2\pi\left(1-\sqrt{1-(k_{\cal{L}}/k_{\mathrm{ER}})^{4}}\right)\,{\rm mod}\,2\pi$ (27)
for $k_{\cal{L}}<k_{\mathrm{ER}}$. Similarly, we find
$\gamma^{LR}_{\mathrm{Berry,u}}=2\pi\left(1+\frac{i}{\sqrt{(k_{\cal{L}}/k_{\mathrm{ER}})^{4}-1}}\right)\;{\rm mod}\;2\pi$ (28)
for $k_{\cal{L}}>k_{\mathrm{ER}}$, and
$\gamma^{LR}_{\mathrm{Berry,u}}=-2\pi\left(1-\frac{1}{\sqrt{1-(k_{\cal{L}}/k_{\mathrm{ER}})^{4}}}\right)\;{\rm mod}\;2\pi$ (29)
[…]
Next, we calculate an integral over the Berry curvature that we call a quasi-Chern integral,
$C^{ab}_{\mathrm{n}}=\frac{1}{2\pi}\int\Omega^{ab}_{\mathrm{n}}(\bm{k})d\textbf{S},$ (30)
with the Berry curvature
$\Omega^{ab}_{\mathrm{n}}(\bm{k})=\frac{\partial A^{ab}_{y,n}(\bm{k})}{\partial k_{x}}-\frac{\partial A^{ab}_{x,n}(\bm{k})}{\partial k_{y}}.$ (31)
Here, the area $S$ is a two-dimensional disk in the $k_{x}-k_{y}$ plane centered at $\bm{k}=0$ with radius $k_{\mathrm{D}}$, either smaller or larger than the ER. In the case without nonlinearity and with zero detuning, we have evaluated the Chern integral analytically and numerically, the latter in order to develop reliable numerics for the case with nonlinearity (discussed below). Since we take the domain of the Berry connection to be the entire $k_{x}-k_{y}$ plane (i.e. $\mathbb{R}^{2}$), rather than a compact and closed Brillouin zone (e.g. a torus), the usual proof of the Chern number being an integer (e.g. [95]) does not hold. While a compact $k$-surface can be divided into two surfaces with one common boundary (and no other boundaries), dividing $\mathbb{R}^{2}$ leads to an additional (non-shared) boundary at infinity [107]. We have seen that in the $RR$ case, the Berry connection is independent of $k_{\cal{L}}$ for $k_{\cal{L}}>k_{\mathrm{ER}}$. This implies that the flux through any annulus bounded by radii $k_{1}$ and $k_{2}$ both larger than $k_{\mathrm{ER}}$ must vanish (because the circulation of the two boundaries of the annulus is equal in magnitude and opposite in sign). Indeed, for the $RR$ case, the Berry curvature is identically zero for all $|\bm{k}|>k_{\mathrm{ER}}$. Therefore the $RR$ Berry curvature becomes independent of the disk radius for $k_{\mathrm{D}}>k_{\mathrm{ER}}$. For $k_{\mathrm{D}}<k_{\mathrm{ER}}$, we find
$C^{RR}_{\mathrm{u}}=-1+\sqrt{1-(k_{\mathrm{D}}/k_{\mathrm{ER}})^{4}}$ (32)
which is as expected from Stoke’s theorem. Hence, outside the ER the quasi-Chern integral is
$C^{RR}_{u}=-1$ (33)
independent of $k_{\mathrm{D}}$. This is our motivation for calling this a quasi-Chern integral. However, we note that the $LR$ result is different. In the $LR$ case, the result for $k_{\mathrm{D}}<k_{\mathrm{ER}}$ is, in agreement with the circulation based on the Berry phase
$C^{LR}_{\mathrm{u}}=-1+\frac{1}{\sqrt{1-(k_{\mathrm{D}}/k_{\mathrm{ER}})^{4}}}.$ (34)
This diverges as $k_{\mathrm{D}}$ approaches $k_{\mathrm{ER}}$, and remains infinite for all $k_{\mathrm{D}}>k_{\mathrm{ER}}$. If the Hamiltonian is written in the standard form of $H=\bm{d(k)}\cdot\bm{\sigma}$ (with $\sigma$ being the Pauli matrices), we see that it maps the 2D $\bm{k}$ plane to a 2D $d_{x}$-$d_{y}$ plane with real-valued $d_{x}$ and $d_{y}$, and a $k$-independent complex z-component $d_{z}=\Delta+i\gamma$. In the $RR$ case, the quasi-Chern integral can be thought of describing the topological properties of $\bm{k}$ in that plane, with the caveat that it is not an integer for disks inside the ER and that it depends on the choice of eigenvectors ($RR$ vs $LR$).